OpenStudy (anonymous):

solve using the quadratic equation-imaginary roots x^2-10x+34

OpenStudy (anonymous):

use the quadratic formula

OpenStudy (anonymous):

Step 1 : Simplify x2+10x + 34

OpenStudy (anonymous):

@eHoaX yea iknow i got to the part of \[110 \pm \sqrt{-116}\div 2\]

OpenStudy (anonymous):

sqrt of a negative turns into an i for imaginary then you proceed as normal

OpenStudy (anonymous):

\[(-b+/-\sqrt{(b)^2-4ac})/(2a)\]

OpenStudy (anonymous):

wait using quadratic formula here omg !

OpenStudy (anonymous):

\[\frac{ 10+/-\sqrt{100-136} }{ 2 }\]

OpenStudy (anonymous):

i think either way works @Shay17

OpenStudy (anonymous):

Solving x2+10x+34 = 0 by the Quadratic Formula . According to the Quadratic Formula, x , the solution for Ax2+Bx+C = 0 , where A, B and C are numbers, often called coefficients, is given by : - B ± √ B2-4AC x = ———————— 2A

OpenStudy (anonymous):

\[\frac{ 10+/-\sqrt{-36} }{ 2 }\]

OpenStudy (anonymous):

\[\frac{ 10+/-6i }{ 2 }\]

OpenStudy (anonymous):

A = 1 B = 10 C = 34 B2 - 4AC = 100 - 136 = -36

OpenStudy (anonymous):

\[5+3i, 5-3i\]

OpenStudy (anonymous):

Applying the quadratic formula : -10 ± √ -36 x = —————— 2

OpenStudy (anonymous):

any more help @a$apkillah

OpenStudy (anonymous):

√ -36 = √ 36 • (-1) = √ 36 • √ -1 = ± √ 36 • i Can √ 36 be simplified ?

OpenStudy (anonymous):

Yes so .. √ 36 = √ 2•2•3•3 =2•3•√ 1 = ± 6 • √ 1 = ± 6

OpenStudy (anonymous):

okay im good i think i understand i messed up thats why thanks @Shay17 and @eHoaX

OpenStudy (anonymous):

So now we are looking at x = ( -10 ± 6i ) / 2 x =(-10+√-36)/2=-5+3i= -5.0000+3.0000i or x =(-10-√-36)/2=-5-3i= -5.0000-3.0000i

OpenStudy (anonymous):

Yw ..

OpenStudy (anonymous):

ima give u a medal @Shay17 and then can u give @eHoaX one so its een since u guys both helped

OpenStudy (anonymous):

even*